# Unseen from Seen: Rewriting Observation-Instruction Using Foundation Models for Augmenting Vision-Language Navigation

### v2

@@ I Introduction
To summarize, the main contributions of this work are:
@@ III Method
In this section, we first describe the VLN problem setup (Section III-A) and then delve into the details of our RAM. The overview of RAM is illustrated in Fig. 2. Specifically, given a human-annotated trajectory-instruction pair, RAM first performs Object-Enriched Observation Rewriting (Section III-B), where Vision-Language Models (VLMs) and Large Language Models (LLMs) are introduced to generate object-enriched rewritten observation descriptions, based on which new observations are synthesized through an efficient panorama-to-view strategy via Text-to-Image Generation Models (T2IMs). Then, RAM conducts Observation-Contrast Instruction Rewriting (Section III-C), which collects grounded landmark information and observation descriptions respectively for the original and new trajectories, and asks the LLMs to contrast the difference between them to produce rewritten instructions. After collecting rewritten observation-instruction pairs, a two-stage mixing-then-focusing training strategy with a random observation cropping scheme is developed to effectively combine the rewritten data with the original data for training (Section III-D). A notation summarization of variables appeared in this section is provided in Table I.
TABLE I: Notation summarization of variables.
Variables Notations
instruction II
observation $O_{t}$
scene description $C_{t}$
rewritten object-enriched scene description $C_{t}^{r}$
rewritten observation $O_{t}^{r}$
ground-truth action (observation) $G_{t}$
grounded landmark for the ground-truth action (observation) $U_{t}$
scene description of the rewritten observation $C^{{}^{\prime}}_{t}$
rewritten instruction $I^{r}$
scene description rewriting prompt $P^{c}$
instruction rewriting prompt $P^{i}$
@@ IV-A2 Evaluation Metrics
TABLE II: Comparison with existing methods on R2R. *: 900 extra simulator scenes for training and using CLIP ViT-H/14 as image features. †: using CLIP ViT-B/16 as image features. ‡: using CLIP ViT-L/14 as image features.
@@
TABLE III: Navigation and object grounding performance on REVERIE. *: 900 extra simulator scenes for training. ‡: using CLIP ViT L/14 as image features.
@@ IV-B Comparison with Existing Approaches
R2R. Table II shows the performance comparison of different methods on R2R, where we can find that our RAM using CLIP ViT-L/14 features outperforms all previous works that do not introduce large-scale additional data for training like ScaleVLN [10]. Although ScaleVLN outperforms previous works as well as RAM, it resorts to significantly larger realistic data resources ($\times$352 larger than the R2R dataset) while our RAM only introduces a small scale of generated data ($\times$3 larger than the R2R dataset). Moreover, our RAM with both two kinds of CLIP features outperforms ScaleVLN [10] in Val Seen split. We can also observe that our RAM outperforms baseline DUET [39] especially in Val Unseen when using both CLIP ViT-B/16 and ViT-L/14 as image features, revealing its superior generalization ability.
REVERIE. Table III compares different methods regarding the navigation and object grounding performance on REVERIE, where we can find that RAM surpasses the baseline [39] largely on both navigation and object grounding metrics under the same visual features on Val Unseen, e.g., the improvements for SR and RGS are $\sim$3.1% and $\sim$2.2%, respectively. RAM also significantly outperforms previous works including AutoVLN [9] and ScaleVLN [10] that introduce large-scale additional simulator data on Test Unseen regarding both navigation and object grounding metrics. These results show that our RAM effectively improves the generalization without the need to introduce large-scale additional realistic data.
TABLE IV: Performance comparison on R4R dataset. ‡: using CLIP ViT-L/14 as image features.
@@
TABLE V: Results on R2R-CE dataset. $\dagger$}: using the same waypoint predictor [64] for a fair comparison.
@@
R4R. Table IV presents the results on R4R, where RAM outperforms the baseline [39] in both navigation (NE, SR, and SPL) and instruction following (nDTW, sDTW, and CLS) metrics on Val Unseen, showing the effectiveness of our method when facing much longer instructions and trajectories. It is also superior to a recent data augmentation method [58] that utilizes T2IMs for environment generation based on the original scene description and adopts Speaker-based instruction augmentation. This comparison result shows that our object-enriched observation rewriting strategy can better improve the scene diversity by introducing new objects and spatial layouts. The comparison also demonstrates the superiority of our observation-contrast instruction rewriting scheme over Speaker-based instruction augmentation.
Transfer to R2R-CE. We further transfer our RAM to R2R-CE. As shown in Table V, although our work is mainly established based on discrete environments, it still shows impressive superiority over the baseline method [39] on R2R-CE, revealing its potential to generalize to continuous environments which is more realistic.
TABLE VI: Ablation results for observation-instruction rewriting. “Pan” stands for generated panoramic images, “Des.” means rewritten object-enriched scene descriptions, and “Ins.” means rewritten instructions, respectively.
@@
TABLE VII: Ablation results of adding data augmented via RAM in different training phases on R2R.
@@ IV-C Ablation Study
We conduct ablation studies to analyze the effects of different method components and the impacts of adding our augmentation data during different training phases. We use CLIP ViT-B/16 features for the ablation experiments. For the baseline in Table II, Fig. 4, and Table VII, we reproduce the baseline model DUET [39] with CLIP VIT-B/16 visual features. In Table VI, since we aim to verify the effectiveness of our rewritten data and exclude the effect of high-quality human annotated instructions, we train the baseline using the widely used R2R augmentation dataset PREVALENT [78], which can provide a fair comparison of different kinds of augmentation data.
@@
TABLE VIII: Comparison with the baseline method [39] under the low-resource settings. Using CLIP ViT-B/16 as image features.
@@
Pretraining & Finetuning Ablation Results. Table VII gives the ablation results of adding RAM augmentation data in different training phases. For No.1 in Table VII, we directly mix the original data with our rewritten data in the pretraining phase, and then finetune the model using only the original data. For No.2 in Table VII, we conduct the pretraining following the baseline model using original data, and then implement our mixing-then-focusing training mechanism using both original and rewritten data in the finetuning phase. In Table VII, through the comparison between the baseline and No.2, we can find that adding our RAM data during the finetuning phase can effectively improve the navigation performance in Val Unseen, showing that our augmentation data is helpful for improving the generalization ability of the VLN agent to unseen scenarios. Moreover, No.3 shows that by further adding our augmentation data in the pretraining phase, the metric results under both Val Seen and Val Unseen can be further boosted. These results show the positive impact of our RAM for enhancing training in different training phases.
@@ IV-D Low-Resource Experiments
We further conduct a low-resource experiment to verify how our RAM augmentation data can help mitigate the data scarcity issue. To this end, we only extract partial original data for the training of baseline [39]. Concretely, we randomly extract the data at different proportions and there may be large distributional discrepancies among extracted training subsets at different proportions. For RAM, we add our rewritten data augmented from the corresponding partial original data. From Table VIII we can find that our RAM surpasses the baseline when extracting original data with different proportions. Especially, our RAM achieves comparable performance to the baseline which uses total original training data in Table VII under the 60% setting. An interesting phenomenon we can also find in Table VIII is that the baseline appears to generalize better with 40% data compared to 60%, which may be caused by large distributional discrepancies among extracted training subsets at different ratios. Moreover, it is also reasonable since it is common that the relationship between performance improvements and data scale is not linear, as also be demonstrated in Fig. 4 for our augmentation data scale ablation. These results further show the practicality of our RAM in real-world applications where the manually annotated data is usually limited.
@@ II-A Scale of augmented data
For each dataset (R2R, REVERIE, and R4R), the augmented data are three times larger than the original dataset, shown in Table IX, which is significantly smaller than other methods such as AutoVLN [9] and ScaleVLN [10]. All our augmentation data are utilized in both pretraining and finetuning phases.
TABLE IX: Scale of augmented data for our RAM and ScaleVLN [10].
@@ II-B Choices of foundation models and implementation details
TABLE X: Details on the time, computational cost, and resources required for generating the augmented data by our RAM.
@@ II-C Details on the time, computational cost and resources
Details on the time, computational cost, and resources required for generating the augmented data by our RAM are presented in Table X. Concretely, we utilized VLM and T2I models to obtain observation descriptions and augmented observations on 8 NVIDIA RTX 3090 GPUs by splitting data into multiple batches in 1 hour and 30 hours, respectively. Rewritten observation descriptions and rewritten instructions were collected by querying LLMs within 30 minutes. Overall, the entire data augmentation pipeline requires less than $\sim$2 days and relatively low computational resources.
TABLE XI: Comparison results with the baseline HAMT [38] on R2R dataset. †: using CLIP ViT-B/16 as image features.
@@ III-A Performance comparison with baseline HAMT on R2R
We have incorporated a new baseline, HAMT [38], along with corresponding experiments on the R2R dataset, as presented in Table XI here. From Table XI, we can observe that our RAM outperforms the baseline HAMT in both Val Seen and Val Unseen, e.g., it brings $\sim$3% improvements in SPL in Val Unseen. The results on different baselines demonstrate the effectiveness and generalization ability of our proposed method.
@@ III-C Comparison with object deletion
To realize object deletion, we mask some specific objects in the observations following EnvEdit [57] and remove the corresponding objects from the rewritten instructions. We present the results here in Table XIII, where we can find that object deletion can also improve performance, indicating that this approach is another effective way to generate meaningful augmentation data. It provides a valuable reference for designing future data augmentation approaches. Furthermore, the results in Table XIII show that our RAM outperforms the object deletion, suggesting that introducing a greater variety of objects can potentially enhance observation diversity and facilitate the agent’s learning of cross-modal alignment, thereby improving navigation performance.
TABLE XIII: Ablation results on object addition and object deletion. We conduct this experiment using CLIP ViT-B/16 as visual features.

### v3

@@ Abstract
Nomenclature
II
Instruction.
$O_{t}$
Observation.
$C_{t}$
Scene description.
$C^{r}_{t}$
Rewritten object-enriched scene description.
$O_{t}^{r}$
Rewritten observation.
$G_{t}$
Ground-truth action (observation).
$U_{t}$
Grounded landmark for the ground-truth action (observation).
$C^{{}^{\prime}}_{t}$
Scene description of the rewritten observation.
$I^{r}$
Rewritten instruction.
$P_{c}$
Scene description rewriting prompt.
$P_{i}$
Instruction rewriting prompt.
@@ I Introduction
To summarize, the main contributions of this work are follows.
  • 1)
@@ III Method
In this section, we first describe the VLN problem setup (Section III-A) and then delve into the details of our RAM. The overview of RAM is illustrated in Fig. 2. Specifically, given a human-annotated trajectory-instruction pair, RAM first performs Object-Enriched Observation Rewriting (Section III-B), where Vision-Language Models (VLMs) and Large Language Models (LLMs) are introduced to generate object-enriched rewritten observation descriptions, based on which new observations are synthesized through an efficient panorama-to-view strategy via Text-to-Image Generation Models (T2IMs). Then, RAM conducts Observation-Contrast Instruction Rewriting (Section III-C), which collects grounded landmark information and observation descriptions respectively for the original and new trajectories, and asks the LLMs to contrast the difference between them to produce rewritten instructions. After collecting rewritten observation-instruction pairs, a two-stage mixing-then-focusing training strategy with a random observation cropping scheme is developed to effectively combine the rewritten data with the original data for training (Section III-D). A notation summarization of variables appeared in this section is provided in Nomenclature.
@@ IV-A2 Evaluation Metrics
TABLE I: Comparison with existing methods on R2R. *: 900 extra simulator scenes for training and using CLIP ViT-H/14 as image features. †: using CLIP ViT-B/16 as image features. ‡: using CLIP ViT-L/14 as image features.
@@
TABLE II: Navigation and object grounding performance on REVERIE. *: 900 extra simulator scenes for training. ‡: using CLIP ViT L/14 as image features.
@@ IV-B Comparison with Existing Approaches
R2R. Table I shows the performance comparison of different methods on R2R, where we can find that our RAM using CLIP ViT-L/14 features outperforms all previous works that do not introduce large-scale additional data for training like ScaleVLN [10]. Although ScaleVLN outperforms previous works as well as RAM, it resorts to significantly larger realistic data resources ($\times$352 larger than the R2R dataset) while our RAM only introduces a small scale of generated data ($\times$3 larger than the R2R dataset). Moreover, our RAM with both two kinds of CLIP features outperforms ScaleVLN [10] in Val Seen split. We can also observe that our RAM outperforms baseline DUET [38] especially in Val Unseen when using both CLIP ViT-B/16 and ViT-L/14 as image features, revealing its superior generalization ability.
REVERIE. Table II compares different methods regarding the navigation and object grounding performance on REVERIE, where we can find that RAM surpasses the baseline [38] largely on both navigation and object grounding metrics under the same visual features on Val Unseen, e.g., the improvements for SR and RGS are $\sim$3.1% and $\sim$2.2%, respectively. RAM also significantly outperforms previous works including AutoVLN [9] and ScaleVLN [10] that introduce large-scale additional simulator data on Test Unseen regarding both navigation and object grounding metrics. These results show that our RAM effectively improves the generalization without the need to introduce large-scale additional realistic data.
TABLE III: Performance comparison on R4R dataset. ‡: using CLIP ViT-L/14 as image features.
@@
TABLE IV: Results on R2R-CE dataset. $\dagger$}: using the same waypoint predictor [63] for a fair comparison.
@@
R4R. Table III presents the results on R4R, where RAM outperforms the baseline [38] in both navigation (NE, SR, and SPL) and instruction following (nDTW, sDTW, and CLS) metrics on Val Unseen, showing the effectiveness of our method when facing much longer instructions and trajectories. It is also superior to a recent data augmentation method [57] that utilizes T2IMs for environment generation based on the original scene description and adopts Speaker-based instruction augmentation. This comparison result shows that our object-enriched observation rewriting strategy can better improve the scene diversity by introducing new objects and spatial layouts. The comparison also demonstrates the superiority of our observation-contrast instruction rewriting scheme over Speaker-based instruction augmentation.
Transfer to R2R-CE. We further transfer our RAM to R2R-CE. As shown in Table IV, although our work is mainly established based on discrete environments, it still shows impressive superiority over the baseline method [38] on R2R-CE, revealing its potential to generalize to continuous environments which is more realistic.
TABLE V: Ablation results for observation-instruction rewriting. “Pan” stands for generated panoramic images, “Des.” means rewritten object-enriched scene descriptions, and “Ins.” means rewritten instructions, respectively.
@@
TABLE VI: Ablation results of adding data augmented via RAM in different training phases on R2R.
@@ IV-C Ablation Study
We conduct ablation studies to analyze the effects of different method components and the impacts of adding our augmentation data during different training phases. We use CLIP ViT-B/16 features for the ablation experiments. In Table V, since we aim to verify the effectiveness of our rewritten data and exclude the effect of high-quality human annotated instructions, we train the baseline using the widely used R2R augmentation dataset PREVALENT [35], which can provide a fair comparison of different kinds of augmentation data. For the baseline in Table I, Fig. 4, and Table VI, we reproduce the baseline model DUET [38] with CLIP VIT-B/16 visual features.
@@
TABLE VII: Comparison with the baseline method [38] under the low-resource settings. Using CLIP ViT-B/16 as image features.
@@
Pretraining & Finetuning Ablation Results. Table VI gives the ablation results of adding RAM augmentation data in different training phases. For No.1 in Table VI, we directly mix the original data with our rewritten data in the pretraining phase, and then finetune the model using only the original data. For No.2 in Table VI, we conduct the pretraining following the baseline model using original data, and then implement our mixing-then-focusing training mechanism using both original and rewritten data in the finetuning phase. In Table VI, through the comparison between the baseline and No.2, we can find that adding our RAM data during the finetuning phase can effectively improve the navigation performance in Val Unseen, showing that our augmentation data is helpful for improving the generalization ability of the VLN agent to unseen scenarios. Moreover, No.3 shows that by further adding our augmentation data in the pretraining phase, the metric results under both Val Seen and Val Unseen can be further boosted. These results show the positive impact of our RAM for enhancing training in different training phases.
@@ IV-D Low-Resource Experiments
We further conduct a low-resource experiment to verify how our RAM augmentation data can help mitigate the data scarcity issue. To this end, we only extract partial original data for the training of baseline [38]. Concretely, we randomly extract the data at different proportions and there may be large distributional discrepancies among extracted training subsets at different proportions. For RAM, we add our rewritten data augmented from the corresponding partial original data. From Table VII we can find that our RAM surpasses the baseline when extracting original data with different proportions. Especially, our RAM achieves comparable performance to the baseline which uses total original training data in Table VI under the 60% setting. An interesting phenomenon we can also find in Table VII is that the baseline appears to generalize better with 40% data compared to 60%, which may be caused by large distributional discrepancies among extracted training subsets at different ratios. Moreover, it is also reasonable since it is common that the relationship between performance improvements and data scale is not linear, as also be demonstrated in Fig. 4 for our augmentation data scale ablation. These results further show the practicality of our RAM in real-world applications where the manually annotated data is usually limited.
@@ II-A Scale of augmented data
For each dataset (R2R, REVERIE, and R4R), the augmented data are three times larger than the original dataset, shown in Table VIII, which is significantly smaller than other methods such as AutoVLN [9] and ScaleVLN [10]. All our augmentation data are utilized in both pretraining and finetuning phases.
TABLE VIII: Scale of augmented data for our RAM and ScaleVLN [10].
@@ II-B Choices of foundation models and implementation details
TABLE IX: Details on the time, computational cost, and resources required for generating the augmented data by our RAM.
@@ II-C Details on the time, computational cost and resources
Details on the time, computational cost, and resources required for generating the augmented data by our RAM are presented in Table IX. Concretely, we utilized VLM and T2I models to obtain observation descriptions and augmented observations on 8 NVIDIA RTX 3090 GPUs by splitting data into multiple batches in 1 hour and 30 hours, respectively. Rewritten observation descriptions and rewritten instructions were collected by querying LLMs within 30 minutes. Overall, the entire data augmentation pipeline requires less than $\sim$2 days and relatively low computational resources.
TABLE X: Comparison results with the baseline HAMT [37] on R2R dataset. †: using CLIP ViT-B/16 as image features.
@@ III-A Performance comparison with baseline HAMT on R2R
We have incorporated a new baseline, HAMT [37], along with corresponding experiments on the R2R dataset, as presented in Table X here. From Table X, we can observe that our RAM outperforms the baseline HAMT in both Val Seen and Val Unseen, e.g., it brings $\sim$3% improvements in SPL in Val Unseen. The results on different baselines demonstrate the effectiveness and generalization ability of our proposed method.
@@ III-C Comparison with object deletion
To realize object deletion, we mask some specific objects in the observations following EnvEdit [56] and remove the corresponding objects from the rewritten instructions. We present the results here in Table XII, where we can find that object deletion can also improve performance, indicating that this approach is another effective way to generate meaningful augmentation data. It provides a valuable reference for designing future data augmentation approaches. Furthermore, the results in Table XII show that our RAM outperforms the object deletion, suggesting that introducing a greater variety of objects can potentially enhance observation diversity and facilitate the agent’s learning of cross-modal alignment, thereby improving navigation performance.
TABLE XII: Ablation results on object addition and object deletion. We conduct this experiment using CLIP ViT-B/16 as visual features.